For a large population of normally distributed account balances, the mean balance is µ  = 150.00 with standard deviation = $35.00. 

(i) What is the probability that one randomly sampled account has a balance that ​exceeds $160.00?  


(ii) What is the probability that the mean for a random sample of n = 40 accounts will exceed $160.00? 


1
Expert's answer
2021-07-26T12:08:25-0400

(i) "X\sim N(150, 35^2)"


"P(X>160)=1-P(X\leq160)"

"=1-P(X\leq \dfrac{160-150}{35})\approx0.387548"


(ii) "\bar{X}\sim N(150, 35^2/40)"


"P(\bar{X}>160)=1-P(\bar{X}\leq160)"

"=1-P(X\leq \dfrac{160-150}{35/\sqrt{40}})\approx0.035380"